If 196g of water is at 38oC, and a copper block at 118oC is added to it, and the final temperature of the mixture is 62oC, what is the mass of the copper?


1
Expert's answer
2021-01-21T03:46:22-0500

Q= mc∆T.

Q= 196 x 4.18 x ( 62-38) = 19662.72

m= Q/c∆T = 19662.72/0.39 x (118-62) = 900,307693.